Grade 12th passPhysical Chemistry

When during electrolysis of a solution of AgN03 9650 coulombs of charge pass through the electroplating bath, the mass of silver deposited on the cathode will be
(a) 10.8 g
(b) 21.6 g
(c) 108 g
(d) 1.08 g

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer ::

No. of moles of Ag = 9650/96500 = 0.1
Mass of Ag deposited = 0.1× 108 = 10.8
Hence option (a) is the answer.


Thanks